if r satisfies the equation r x (i +2 j+ k)= i-k then for any scalar m , r is eqaul to?????

Grade:12th pass

6 Answers

ng29
8 years ago
consider  r =ai+bj+ck
and solve lhs by using crosss product now  get single equations in i,j and k
compare coefficients of i, j and k and u get the desired vector
 
approve if useful
ng29
8 years ago
wait …..wait........
 
what is m ?????
mansi dabriwal
8 years ago
i tried through that only ….but don know what this m...stands for...
options for this ques. …
a) i+m(i+2j+k)
​b) j+ m(i+2j+k)
​c) k+m(i+2j+k)
​d) i- k +m (i+2j+k)
ng29
8 years ago
answer is b)
now i understand that what is m 
when u compare the both sides u does not get exact coeffiicients of r 
i am solving completely for u
let r=ai+bj+ck
on taking cross product and comparing we get   a=c ,  b=2c+1
put these values in r
we get r = ci+(2c+1)j+ck
 r = j + c(i+2j+k)
in options they had considered m in place of c
 
APPROVE IF USEFUL
mansi dabriwal
8 years ago
k...i got it...thanx...
ng29
8 years ago
my pleasure